Parth (parthkohli):

If \(x^{100} + 1 \) is divided by \(x^2 - 1\), then what is the remainder?

Parth (parthkohli):

Polynomial remainder theorem doesn't work here, unfortunately. Does it?

Parth (parthkohli):

\( \color{Black}{\Rightarrow x^2 - 1 = (x + 1)(x - 1) }\)

OpenStudy (anonymous):

this may help u http://in.answers.yahoo.com/question/index?qid=20120328050459AASixcw

Parth (parthkohli):

I need to work on my own, and I need step-by-step guidance.

mathslover (mathslover):

x^2-1=(x+1)(x-1) .. first step:)

mathslover (mathslover):

i got something please wait

mathslover (mathslover):

x^2-1=0 x^2=1 x=1 x^(100)+1=1^(100)+1=2

Parth (parthkohli):

I see.

mathslover (mathslover):

remainder theorem says that .. if p(x) is divided by p(a) then

mathslover (mathslover):

sorry leave that : let x^2-1=0

mathslover (mathslover):

then x = ?

Parth (parthkohli):

How can we even divide a number by 0?

mathslover (mathslover):

no dont divide .. equate that as 0 \[\large{x^2-1=0}\] x =?

Parth (parthkohli):

\(x = \pm 1\)

Parth (parthkohli):

Now we can find \(p(1)\)

mathslover (mathslover):

right now put that in p(x) = \(\large{x^{100}+1}\) ..p(1)=? remainder

Parth (parthkohli):

I do know the remainder theorem for \(p(x) \div x - k\), but I didn't know how to apply that here. Why do we equate with 0?

mathslover (mathslover):

we generally : suppose x-k=0 that is x =k and then remainder = p(k) ..

Parth (parthkohli):

Oh, I see.

Parth (parthkohli):

Thank you!

mathslover (mathslover):

did that seriously helped you .. i dont think i did

mathslover (mathslover):

I think you needed the proof for that thingy ?

Parth (parthkohli):

No, I understand it fully now! I do have it.

mathslover (mathslover):

ok just for reference : ncert IX chapter 2 has explanation for this theorem

mathslover (mathslover):

well nice to hear that u got it .. :)

Parth (parthkohli):

I do have it :P

mathslover (mathslover):

:P good

Parth (parthkohli):

Thanks again!

mathslover (mathslover):

no problem.. have to go now bbye

Parth (parthkohli):

Bye! :)